Solve the system by graphing, addition and substitution methods.
–4x + 7y = –10
x – 5y = 9

Answer by checkley71(8403)

-4X+7Y=-10 NOW MULTIPLY EQ #2 BY 4 WE GET
4X-20Y=36 NOW WE ADD THESE 2 EQ.
------------------
-13Y=26
Y=26/-13
Y=-2
-4X+7*-2=-10
-4X-14=-10
-4X=14-10
-4X=4
X=4/-4
X=-1
PROOF
-1-5*-2=9
-1+10=9
9=9